# Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v3

@@ 2.3.2 Our contribution: Extracting the in-context prior and posterior
These are distributions that estimate $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$. For a more detailed breakdown of the process, see Appendix A.2. In our experiments, to sample each $\varphi_{\text{MLE}}$, we choose to sample $X\sim\mathcal{U}[-5,5]$ five times for each of the $100$ task descriptions. Since we normalise our real data, these samples of $X$ will cover the domain of $f_{\varphi}$ with a high probability.
In the linear case (for both linear and binary logistic regression), this process is simple since we can calculate the maximum likelihood estimate of $y=\varphi_{0}+\varphi\cdot x$ in closed form since we have direct access to $p_{M,T}(y|x,\varphi)$, which is either a regression label or the logit of a class.
Since we now have access to estimates of both $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$, we can investigate whether the language model is accurately conveying it’s internal prior when it provides us with a prior distribution through elicitation and, when doing in-context learning, whether it is performing Bayesian inference.
2.3.3 Our contribution: Is the language model using the elicited prior in-context?
Since we now have access to an approximation of the prior on a language model’s own internal predictive model $p_{M,T}(\varphi)$, we can study how “truthful” the language model is when it provides us with a prior for a separate model $p_{M,T}(\theta)$ through elicitation (as in Section 2.2 and 3.1). By studying these differences, we can understand one of the trade-offs between using in-context learning and prior elicitation; the language model may not be able to reliably convey its own prior knowledge of a predictive task when we elicit a distribution, or is unable to translate its own prior knowledge into prior distributions over the parameters of its internal predictive model. Either case leads to differences in the predictive models.
To quantify this, we use the energy statistic [33] (Appendix A.4), which measures how closely two distributions match (taking values $\in[0,1]$ with 00 indicating that two distributions are equal). Here, we test the difference between the parameter distributions from either the elicited distribution ($\theta\sim p_{M,T}(\theta)$) or the one approximated from the language model’s in-context predictive model ($\varphi\sim p_{M,T}(\varphi)$).
2.3.4 Our contribution: Inspecting the in-context posterior distribution
As part of the decision between using in-context learning and prior elicitation, we would like to understand whether the language model is performing Bayesian inference when it makes predictions in-context.
We can use the approximated internal prior $p_{M,T}(\varphi)$ along with Monte Carlo methods [34, 35] (Appendix A.8) to sample the posterior, and compare it to the approximated internal model posterior $p_{M,T}(\varphi|D)$ presented in Figure 5 and extracted using the MLE sampling introduced in Section 3.3. The difference in these distributions can be measured using the energy statistic as in Section 3.4 to understand whether the language model is acting predictably after observing data from $\mathcal{D}$.
If the language model’s internal posterior is significantly different from the one calculated by applying Monte Carlo methods to the approximated prior, then in-context learning might be considered unusable for a given task, as it is not clear that the model is updating its prior belief about a predictive task from the data. This would suggest that the language model should not be used to generate posterior distributions on predictions in-context as it is improperly applying Bayes rule after observing demonstrations.
@@ 2.4 Bayes factor for model selection
To decide whether to use in-context learning or prior elicitation, given a data set, we consider the Bayes factor for model selection.
The Bayes factor compares the marginal likelihood ratio of two statistical models to determine under which model the data was most likely generated. Formally, given two models $\alpha_{0}$ and $\alpha_{1}$ with parameterisations $\theta_{0}$ and $\theta_{1}$, the Bayes factor calculated over a dataset $\mathcal{D}$ is:
[…]
In our case, $\alpha_{0}$ refers to prior elicitation, with the prior determining the parameterisation $\theta$ of a predictive model, and $\alpha_{1}$ referring to in-context learning. We can approximate the Bayes factor using the means from samples of the likelihood of $D$ (Section 3.6).
This method of measurement is only applicable if the language model is performing Bayesian inference when learning in-context, which is explored in Section 3.5.
2.5 Experimental datasets
Within our experiments, we present results on five real-world datasets and a synthetic dataset, covering both classification and regression tasks from various domains.
The equation $y=2x_{1}-x_{2}+x_{3}$ is used to generate a synthetic dataset that contains 3 features sampled from a normal distribution (mean $=0$ and standard deviation $=1$) and targets calculated using the equation and perturbed by noise $\epsilon\sim\mathcal{N}(0,0.05)$. When describing this dataset to the language model, we fully specify the relationship between the features and targets to test the setting where the task description is fully descriptive.
To demonstrate the effectiveness of the LLM elicited priors on new tasks, we use a private dataset (i.e., not seen by the LLM) collected as part of a study on Dementia care which contains $10$ features of in-home activity and physiology and labels corresponding to days with a clinically validated labelling of positive or negative urinary tract infection (UTI), calculated through the analysis of a provided urine sample in consultation with a healthcare professional. For this analysis, we used a subset of data containing $78$ people and $371$ positive or negative days of UTI. This dataset has not appeared online and will not have been part of the language model’s training.
We then test four public benchmark datasets. The Heart Disease dataset [36, 37] where the task is to predict a heart disease diagnosis from physiological features; the Diabetes dataset [38], with the goal of predicting a quantitative measure of disease progression one year after baseline; the Hypothyroid dataset [39, 40], where the task is to predict whether a patient has hypothyroidism based on biomarkers; and the Breast Cancer dataset [41], where the task is to predict the diagnosis of a breast mass.
Further information on the contents of these datasets is given in Appendix A.5.1. In Appendix A.9, we apply two of the memorisation tests proposed in Bordt et al. [42] to explore the extent to which the language model has memorised the public datasets. The synthetic and UTI datasets are generated or private, and therefore are unseen to the language model.
@@ 3 Results
All experiments in this section were conducted with OpenAI’s GPT-3.5-turbo, with further language models tested in Appendices A.13.1 and A.13.2.
3.1 Evaluating the LLM-elicited priors
Figure 2: Test accuracy of the posterior distribution for varied numbers of training data points. Each of the Gaussians in the mixture is extracted from a language model and is tested against an uninformative prior $\mathcal{N}(\theta|0,1)$ on a linear model. These results are calculated on a test set of the datasets after each model is trained on the given number of data points. These values are the result of $5000$ sampled parameters and $10$ random splits of the training and testing data. For Diabetes, a regression task, we report the mean squared error on log-scale. 2(a) shows the histogram of accuracy achieved whilst 2(b) shows a line plot of the average mean posterior accuracy or Mean Squared Error (MSE), with the error bars representing the 95% confidence interval, as well as the additional datasets tested in Appendix A.6. The green arrows point in the direction of improvement in the metric.
We start by exploring the performance difference between a predictive model using an uninformative prior distribution over our parameters, $\theta\sim\mathcal{N}(\theta|0,1)$, with one elicited from a language model. For each dataset, we perform cross-validation with five splits and calculate the posterior performance as we vary the number of training data points.
Figure 2(a) shows the results of these experiments, demonstrating that the prior distribution on the linear model’s weights elicited from a language model provides better mean accuracy than an uninformative prior for smaller numbers of data points and often provides less variation in accuracy.
On the synthetic data, we learn to model $y=2x_{1}-x_{2}+x_{3}$ using a prior elicited from the language model. In this case, since we are testing how well the language model can provide priors based on informative task descriptions, we describe the exact relationships between $x$ and $y$ in the task descriptions. The LLM-elicited prior provides a lower MSE, by an order of magnitude, compared to an uninformative prior on the test set after $5$ observed examples and an equivalent MSE after $20$ data points are seen. This is likely because the task was simple enough to be learnt effectively by an uninformative prior after a few examples. However, it demonstrates that the language model is capable of providing effective prior distributions given a clear task description. In fact, the mixture of priors that the language suggests are shown in Appendix A.6.1 and are centred on the true parameter values, suggesting that the language model is able to convert its combined knowledge of a problem and a task description into useful prior distributions. The relationship between the details of the descriptions and the performance of elicited priors is studied within Appendix A.10.
Figure 2(b) shows that for real-world datasets, we often see the same posterior performance improvements. On the Breast Cancer dataset, the improvement in accuracy is clear up to $40$ observed data points, where the mean accuracy is greater, and its variation is smaller. The elicited prior distribution achieved a greatly increased accuracy for lower numbers of observed data points, enabling more reliable predictions earlier into model development.
In the case of Hypothyroid prediction, using the elicited priors did not improve the posterior predictions for low numbers of observed data points. However, once the number of observed data points increased to $20$ and greater, the accuracy was higher when using the elicited priors. As in Section 2.4 and 3.6, with a Bayesian approach, the likelihood of the prior distributions can be measured given the observed data before they are used for new predictions and the appropriate prior can be selected without access to a test set.
Interestingly, for Heart Disease, there is no noticeable difference between the average accuracy of a logistic regression model under an uninformative or elicited prior. This is because the elicited prior contains many Gaussian distributions of mean 00 and standard deviation $1$, making it very similar to the uninformative prior. This is illustrated in Figure 12 within Appendix A.7 which plots the elicited parameter distributions as a histogram in each feature dimension, for the first $10$ features. This is reassuring, since it shows that the language model is able to respond with an uninformative prior when it is not confident in its own prior knowledge of a task.
For both the Diabetes and UTI datasets, the elicited prior provides a greater mean accuracy or lower MSE over the uninformative prior for all of the number of observed data points tested. This should be of particular note considering that the UTI dataset is private and will have never appeared within the language model’s training.
In general, we see that the greatest improvement in accuracy occurs for smaller numbers of observed samples, suggesting that the use of prior knowledge from language models is most relevant when users have access to only a few samples. This can be important in scenarios where collecting labels is costly and time-intensive, as is the case in the UTI dataset. In Appendix A.6.2, we consider the accuracy achieved using uninformative priors and elicited priors in the context of collection time. We show here that a UTI predictive model might have achieved greater accuracy earlier in the study. In a clinical environment, this could have enabled improved UTI detection earlier, accelerating model development and validation.
In Appendix A.6, we present additional accuracy distributions (similar to Figure 2(a)), using the Heart Disease and Hypothyroid datasets, which supplement the results in Figure 2(b). In Appendix A.11, we examine the number of task descriptions required to elicit consistent prior distributions, and how the detail of these descriptions affects posterior performance in A.10.
3.2 Incorporating expert information
Figure 3: UTI prior parameters with expert information. The histograms, in the three left hand plots, show the distribution of parameter values for the features that we provided expert information for. In all three cases, we stated that the three features are positively correlated with UTI risk. In the right hand plot, we show the posterior accuracy of these prior distributions for different numbers of observed samples. The line shows the mean posterior accuracy and the error bars represent the 95% confidence interval over the $10$ test splits.
By including expert information in the task description, acquired through conversations with domain experts, we can steer a language model to provide prior distributions that combine knowledge from clinicians with its own learnt insights.
@@ 3.2 Incorporating expert information
To study this, we chose to also include our own knowledge of the risk factors for UTI in the task description for prior elicitation [43]. In particular, we provide further information on three features available in the dataset that are positively correlated with the outcome. We added the following phrasing to the system prompts, based on clinical knowledge of UTIs: “It is known that the more previous UTIs a patient had, the more likely they are to have a UTI in the future. Someone with a UTI will also wake up more during the night. Someone with a UTI will have a greater bathroom usage.”
Figure 3 shows the effect on the prior parameter distribution this has, where all three features we mention in the prompt become more positively predictive of a UTI.
It should be noted that without this expert information, the LLM had already correctly identified the relationship between the risk of UTI and bathroom usage, the frequency of nighttime disturbances, and the number of previous UTIs, as these are all positive predictors of UTI. The language model had already translated this information from its internal prior into the elicited distributions. This could explain the small improvement in posterior accuracy over LLM-elicited priors without additional expert information, shown in the line graph of Figure 3.
In Appendix A.10, we examine the elicited priors when incrementally including more detailed expert information in the task description for a synthetic dataset, which a language model holds no prior knowledge for.
3.3 Extracting the in-context prior and posterior
@@
Figure 4: Extraction of the language model’s hidden prior for in-context predictions. The bottom row shows the mean squared error between the logits or regression predictions outputted by the language model and the values calculated by the fitted MLE models, whilst the top row shows histograms of the parameter values of these MLE models, with each colour representing a single feature. For ease of visualisation, we exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles and the bias term. In the top row, each colour represents a single feature.
@@
Figure 5: Extraction of the language model’s hidden posterior for in-context predictions. The bottom row shows the mean squared error between the logits or regression predictions outputted by the language model and the values calculated by the fitted MLE models, whilst the top row shows histograms of the parameter values of these MLE models, with each colour representing a single feature. For ease of visualisation, we exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles and the bias term. In the top row, each colour represents a single feature.
This section explores the prior and posterior distributions of the linear predictive model that the LLM is using in-context (Appendix A.5.3).
The prior and posterior distributions estimated through a repeated sampling of the in-context predictions (Section 2.3.2) are shown in Figures 4 and 5. To quantify our approximation error, once we estimate the parameters of the linear model, we can use them to calculate the predictions returned by the language model in-context to measure the difference between the output of the language model and the estimated internal model. This is shown in the bottom row of Figures 4 and 5, suggesting that our estimates of the internal parameters are reliable for the real-world datasets.
For the prior distributions in Figure 4, we see variations in the estimated prior distributions across the datasets, suggesting that the language model uses some knowledge about the tasks when making predictions, even without any demonstration examples. However, we see that the language model often uses a prior centred close to 00, suggesting that it is not particularly informative.
The extracted posterior distributions are presented in Figure 5 and show some differences compared to the prior distributions. For the Diabetes and Heart Disease tasks our estimation error increased but remained low. However, we find that the Breast Cancer and Hypothyroid dataset produces a large error for our predictions of the logits (Figure 5), suggesting that our estimates of the internal posterior model parameters are not reliable for this case. The error in approximating the parameters of the in-context model for the synthetic task is discussed in Appendix A.3. It is interesting that, on the whole, we were able to produce a better approximation of the language model’s internal prior distribution than the posterior distribution. This is likely due to the limitations of the language model, and in particular, its capacity to understand demonstrations and instructions in the same prompt.
As LLMs improve their ability to follow instructions, these approximation errors will reduce considerably since they are a product of the language model’s inconsistent imitation of a linear model.
We do not estimate the language model’s posterior distribution for the UTI dataset as samples in this dataset are kept private, and so we are unable to provide demonstrations in-context.
3.4 Is the language model being truthful?
Table 1: Extracted prior compared to the elicited prior. The energy statistic between the language model’s internal prior extracted through MLE sampling and the prior elicited by following the method in Section 2.2. Values are $\in[0,1]$ with values closer to zero, indicating that the two distributions are more similar.

### v4

@@ 4.2 In-context learning
In our experiments, for each $\varphi_{\text{MLE}}$, we sample 25 data points from $X\sim\mathcal{U}[-5,5]$, five times for each of the $100$ task descriptions. As the real data is normalised, these samples of $X$ will cover the domain of $f_{\varphi}$ with a high probability. Appendix A.3 provides more details on the process.
When specifying linear $f_{\varphi}$, we calculate the MLE in closed form as we have $y=\varphi_{0}+\varphi\cdot x$, which is either a regression label or the logit of a class. Therefore, large approximation errors will be due to the LLM not using a linear $f_{\varphi}$.
By repeated sampling, we form distributions that estimate $p_{M,T}(\varphi)$ and $p_{M,T}(\varphi|\mathcal{D})$. These are over the hidden parameters of the in-context model and give us new visibility into the LLM’s predictive process; assuming the model class can be specified in a prompt (Appendix A.4). Using these, we can test whether an LLM is using the elicited prior in-context and if it is approximating Bayesian inference.
Is the LLM using the elicited prior in-context?
To quantify the difference between the in-context prior ($p_{M,T}(\varphi)$) and the one we elicit ($p_{M,T}(\theta)$), we will use the energy statistic (Székely & Rizzo, 2013). This takes values $\in[0,1]$ with 00 indicating that the two distributions are equal (Appendix A.5) and evaluates how consistently the LLM uses its background knowledge.
The in-context posterior
We also test whether the LLM is approximating Bayesian inference during in-context learning. We use the approximated in-context prior $p_{M,T}(\varphi)$ and Monte Carlo methods (Hastings, 1970; Hoffman et al., 2014) (Appendix A.14) to sample a posterior, and compare it to the extracted in-context posterior $p_{M,T}(\varphi|D)$. The energy statistic quantifies the difference in these distributions. If they differ, it suggests the LLM is not performing Bayesian inference in-context and is not acting predictably after observing data from $\mathcal{D}$.
@@ 4.3 Bayes factor for model selection
If the LLM is approximating Bayesian inference in-context, we use the Bayes factor for model selection between in-context learning and AutoElicit priors. This calculates the marginal likelihood ratio of two statistical models to determine under which the data was likely generated (Appendix A.6), defining a model selection strategy given any dataset.
For AutoElicit (method $\alpha_{0}$), we approximate the marginal likelihood by sampling the prior predictive log-likelihood over $\mathcal{D}$. We sample $500$ predictions for $25$ data points to get samples of $p_{M,T}(\mathcal{D}|\alpha_{0})$. For in-context learning (method $\alpha_{1}$), we approximate $p_{M,T}(\mathcal{D}|\alpha_{1})$ as the LLM provides predictions on features directly. We prompt the LLM with $100$ task descriptions and ask it to predict on $25$ data points from the dataset. This is repeated over $5$ splits of the datasets. Further details are given in Appendix A.7.5.
@@ 5 Results
5.1 Experimental setting
We present results on one synthetic and five clinical datasets, covering classification and regression (Appendix A.7.1).
The synthetic data set contains 3 features from $X\sim\mathcal{N}(0,1)$ and targets calculated using $y=2x_{1}-x_{2}+x_{3}+\epsilon$, with $\epsilon\sim\mathcal{N}(0,0.05)$. When describing this task to the LLM, we fully specify the relationship between the features and targets to control for the completeness of the descriptions.
We test four publicly available datasets. The Heart Disease dataset (Detrano et al., 1989): predicting a heart disease diagnosis from physiological features; the Diabetes dataset (Efron et al., 2004): predicting disease progression one year after baseline; the Hypothyroid dataset (Quinlan, 1986; Turney, 1994): predicting hypothyroidism based on biomarkers; and the Breast Cancer dataset (Wolberg et al., 1995): predicting the diagnosis of a breast mass.
We also evaluate AutoElicit on a private dataset not in the LLM’s training, collected as part of our study on Dementia care (Capstick et al., 2024). This contains $10$ features of daily in-home activity and physiology data, and clinically validated labels of positive or negative urinary tract infection (UTI), based on the analysis of a urine sample by a healthcare professional. This dataset contains $78$ people and $371$ positive or negative days of UTI.
To explore if the LLM has memorised this data, we use two tests (Bordt et al., 2024) in Appendix A.8. Appendix A.7.6 explains how to use these methods on a new dataset.
@@ 5.1 Experimental setting
All experiments in this section were conducted with OpenAI’s GPT-3.5-turbo and GPT-4-turbo, and DeepSeek-R1-32B with 4-bit quantisation. All details are given in Appendix A.7 and other LLMs are tested in Appendix A.15.
5.2 How informative are AutoElicit priors?
We compare the posterior performance of a predictive model using an uninformative prior $\theta\sim\mathcal{N}(\theta|0,1)$, with those elicited from three LLMs. Figure 2 shows that AutoElicit priors provide substantially better accuracy and mean squared error (MSE) than an uninformative prior, and at least one of the LLMs provides the greatest performance for all training sizes and tasks. The complete accuracy distributions and a detailed analysis is given in Appendix A.9; Appendix A.10 describes the associated cost; and A.12 studies how the number of descriptions affects the priors.
Figure 2: Test performance of the posterior distribution for varied training data sizes. The average and 95% confidence interval of the mean posterior accuracy or Mean Squared Error (MSE) of $10$ splits of the dataset using AutoElicit priors and an uninformative prior ($\theta\sim\mathcal{N}(\theta|0,1)$). These are calculated on a test set of each dataset, with the green arrow pointing in the direction of metric improvement.
When eliciting priors for $y=2x_{1}-x_{2}+x_{3}$, we specify the equation in the prompt to test the LLMs’ ability to provide priors given complete descriptions (incomplete descriptions are studied in Appendix A.11). In Figure 2, the LLM-elicited priors provide an order of magnitude lower test MSE compared to an uninformative prior after $5$ observed examples and an equivalent MSE after $20$, as the task was simple enough to be learnt with only a few examples. GPT-3.5-turbo provided the best prior as it returned sharp distributions (Appendix A.9.1), reflecting the true parameter values. This shows LLMs can combine their knowledge of probability with task descriptions to form useful priors.
For Heart Disease, uninformative and GPT-3.5-turbo-elicited priors produced the same mean posterior accuracy, as this elicited prior contains many standard Normal distributions (Figure 15). This is reassuring, as GPT-3.5-turbo responded with an uninformative prior when unsure of its knowledge. However, GPT-4-turbo and DeepSeek-R1-32B-Q4, returned considerably better priors. The appropriate prior can be chosen with Bayes factor (Appendix A.16)
Of particular note are the results on the UTI task, a private dataset where all AutoElicit priors provide greater accuracy over the uninformative prior for all training sizes. Figure 1 shows these results considered alongside label collection time, illustrating that we achieve greater accuracy earlier in the study, accelerating model development and validation. This is discussed further in Appendix A.9.2.
Also of importance are the results on the classification tasks of the priors elicited from DeepSeek-R1-32B-Q4, a 4-bit quantised version of the distilled DeepSeek-R1 model. This suggests the reasoning qualities of more recently released LLMs are of significant use for eliciting priors.
These results illustrate the remarkable effectiveness of using AutoElicit priors, which considerably improve posterior predictions, particularly for smaller training sizes; where label collection might have been challenging. Our framework makes expert prior elicitation more accessible.
5.3 Can experts supplement the process?
@@
Including expert information in the task description allows AutoElicit priors to combine learnt insights with domain knowledge (Capstick et al., 2024). In the UTI task description we now list three features that are positively correlated with the label: “It is known that the more previous UTIs a patient had, the more likely they are to have a UTI in the future. Someone with a UTI will also wake up more during the night. Someone with a UTI will have a greater bathroom usage.”
Figure 3 shows that this information is incorporated when LLMs provide priors, as these three features become more positively predictive of a UTI. GPT-3.5-turbo had already identified this relationship, shown by the positive median of the priors, which explains the small improvement in posterior accuracy this additional information provides. Further, Appendix A.11, studies the priors elicited as we incrementally improve the description of a synthetic task, illustrating that GPT-3.5-turbo produces reliable distributions from text.
Figure 3: UTI prior parameters with expert information. The histograms in the three left hand plots show the distribution of parameter values for the features that we provide expert information for in the task description. In all three cases, we state that the features are positively correlated with UTI risk. In the right hand plot, we show the posterior accuracy of these distributions for different numbers of observed samples. The average and 95% confidence interval of the mean posterior accuracy over the $10$ test splits are shown.
AutoElicit can therefore also be used to supplement typical prior elicitation methods, with the benefit that experts can convey their knowledge as natural language.
5.4 The in-context priors and posteriors
Figures 4 and 5 show GPT-3.5-turbo’s in-context prior and posterior distributions approximated with maximum likelihood estimation (MLE). To quantify the error, we calculate the mean squared error (MSE) between the LLM’s in-context predictions and the approximated model’s predictions, shown in the bottom row of Figures 4 and 5. This allows us to study the implicit parameters used by the LLM to make predictions in-context.
@@ 5.4 The in-context priors and posteriors
Figure 4: Extraction of GPT-3.5-turbo’s in-context prior. The bottom row shows the MSE between the LLM’s in-context predictions and the MLE model’s predictions, whilst the top row shows the parameter distribution of these MLE models, with each colour representing a single feature. For ease of visualisation, we exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles and the bias term.
@@
Figure 5: Extraction of GPT-3.5-turbo’s in-context posterior. The bottom row shows the MSE between the LLM’s in-context predictions and the MLE model’s predictions, whilst the top row shows the parameter distribution of these MLE models, with each colour representing a single feature. For ease of visualisation, we exclude values outside of the $(2.5\%,~{}97.5\%)$ percentiles and the bias term.
The estimated priors in Figure 4 show variations in their values across the datasets, suggesting the LLM uses its own knowledge in-context before observing demonstrations. The low MSE on the real-world tasks indicates that the LLM was reliably imitating a linear model, allowing us to achieve a low approximation error. However, the priors are not particularly informative, as they are often centred close to 00.
The estimated in-context posteriors in Figure 5 differ from the priors, suggesting the LLM has updated its knowledge based on demonstrations. For the Diabetes and Heart Disease tasks, the LLM mostly continued to produce linear predictions, and so our error increases, but remains low. However, the Breast Cancer and Hypothyroid dataset produce a large approximation error, as the LLM struggled to imitate a linear model. The approximation error on the synthetic task is discussed in Appendix A.4. We do not estimate the in-context posterior for the UTI dataset as it is private, and so we are unable to provide demonstrations.
In general, we better approximate the LLM’s in-context prior than posterior. This is likely due to limitations in the LLM’s capacity to interpret demonstrations for training, alongside instructions to provide linear predictions, in the same prompt. As the abilities of LLMs improve, we expect these approximation errors to reduce, as they are due to the language model’s inconsistent imitation of a linear model.
This is an approximation of the LLM’s in-context parameter distribution, giving new visibility into its predictive process.
5.5 Is the language model consistent?
We compare these in-context priors with the AutoElicit priors. Table 1 suggests the LLM uses different priors during in-context learning compared to those we elicit, since they differ according to their energy statistic. This is supported by Appendix A.13, which presents histograms of the parameter values. Either the distribution we elicit is an unreliable representation of the LLM’s knowledge, or it is unable to use this knowledge when making predictions in-context. As in-context priors are often centred around 00, we suspect it is the latter, and that prior elicitation more effectively makes use of the LLM’s background knowledge.
Table 1: Approximated GPT-3.5-turbo in-context prior compared to the elicited prior. The energy statistic between the LLM’s approximate in-context prior, and the LLM-elicited prior. Lower values indicate that the two distributions are more similar.